1. Suppose the value of the price elasticity of demand is -3.Interpret this number if the price increasedby 1%.

2. (3points)Suppose that when the price of beer is $2 per bottle, firms can sell 4 million bottles. When the price of beer is $3 per bottle, firms can sell 2 million bottles.

A. Using the midpoint method, calculate the elasticity between $2 and $3. B.Calculate the difference in total revenue when the price increasesfrom $2 to $3.

B. Calculate the difference in total revenue when the price increases from $2 to $3.
Answer #1

1) Elasticity = -3

so when the price increases by 1%, the quantity demanded will decrease by 1*-3 = -3%

2) E = [2-4/(2+4/2)] / [3-2/(3+2/2)] = [2/3] / [1/2.5] = 0.6666/0.4 = 1.6

b) difference in total revenue

when price is 2, TR = 2*4 =8 million

when price is 3, TR = 3*2 = 6 million

so, TR decreases by 8-6 = 2 million